# RIFT: A RubrIc Failure Mode Taxonomy and Automated Diagnostics

### 01 Apr 2026 v1

@@ Abstract
Rubric-based evaluation is widely used in LLM benchmarks and training pipelines for open-ended, less verifiable tasks. While prior work has demonstrated the effectiveness of rubrics using downstream signals such as reinforcement learning outcomes, there remains no principled way to diagnose rubric quality issues from such aggregated or downstream signals alone. To address this gap, we introduce RIFT: RubrIc Failure mode Taxonomy, a taxonomy for systematically characterizing failure modes in rubric composition and design. RIFT consists of eight failure modes organized into three high-level categories: Reliability Failures, Content Validity Failures, and Consequential Validity Failures. RIFT is developed using grounded theory by iteratively annotating rubrics drawn from five diverse benchmarks spanning general instruction following, code generation, creative writing, and expert-level deep research, until no new failure modes are identified. We evaluate the consistency of the taxonomy by measuring agreement among independent human annotators, observing fair agreement overall (87% pairwise agreement and 0.64 average Cohen’s kappa). Finally, to support scalable diagnosis, we propose automated rubric quality metrics and show that they align with human failure-mode annotations, achieving up to 0.86 F1.
@@ 2 Related Work
Issues With LLM Evaluation Benchmarks. Recent studies show that LLM benchmarks suffer from unreliable outcome verification, including widely used benchmarks such as SWE-bench-Verified (Chowdhury et al., 2024) and $\tau$-bench (Yao et al., 2024) (Zhu et al., 2025b). Beyond outcome verification, naive LLM-as-Judge evaluation can introduce systematic bias and overconfident estimates which requires calibrated reporting (Lee et al., 2026). Prior work also reports substantial evaluation variance (Madaan et al., 2024) and other weaknesses in current benchmarking practices (Reuel et al., 2024; Eriksson et al., 2025), motivating the need for more reliable evaluation design.
@@ 3 The RubrIc Failure mode Taxonomy
Table 1: RIFT: RubrIc Failure Mode Taxonomy. See Appendix C for complete descriptions.
@@
Reliability Failures Subjective Use of inherently subjective evaluative terms without an anchor to objective expectations. 0.59 0.60 86.7%
Non- Atomic Missing a parseable, consistently scorable structure, or presence of bundled criteria that prevents partial credit. 0.54 0.54 86.7%
Ungrounded One or more criteria require verification against information that is factually checkable, but fail to provide sufficient grounding or bounds. 0.73 0.74 86.7%
Content Validity Failures Misaligned or Rigid The rubric (a) grades the wrong objective, or (b) imposes unnecessarily strict requirements. 0.57 0.58 80.0%
Missing Criteria The rubric provides no criterion to evaluate at least one expected requirement. 0.69 0.69 86.7%
Consequential Validity Failures Hackable A responder could easily achieve a top score without materially improving quality. 0.74 0.72 93.3%
Low Signal The rubric would give nearly equivalent scores to many different-quality responses. 0.86 0.86 93.3%
Redundant Criteria Two or more criteria evaluate the same requirement, such that the same behavior is rewarded or penalized multiple times. 0.70 0.67 86.7%
@@
We define the RubrIc Failure mode Taxonomy (RIFT) as a generic framework for identifying and classifying failure modes in evaluation rubrics (Tab. 1). To avoid imposing a top-down schema, RIFT is developed using grounded theory (Glaser and Strauss, 1967a), in which failure modes are iteratively derived from expert annotations and open-ended feedback. The resulting taxonomy organizes rubric failures into three core categories: Reliability (consistency and reproducibility of judgments), Content Validity (alignment between rubric criteria and the intended evaluation target), and Consequential Validity (downstream usefulness and discriminative power of the rubric).
@@
Taxonomy Development Pipeline. We develop RIFT using a constant comparative method over four sequential iterations, with the 85 rubrics split into four subsets and processed in order. Each iteration consists of three individual expert annotations followed by panel taxonomy discussion and revision; the first iteration additionally includes a bootstrapping step to initialize the taxonomy.
@@
    Bootstrapping (first iteration only). Experts independently write open-ended rubric critiques for each rubric; the critiques are provided to GPT-5.2 to propose an initial taxonomy.
@@
    Labeling and feedback. Experts label each rubric using the current taxonomy and document failure modes not captured by existing categories, as well as issues in the taxonomy itself (e.g., description inaccuracies and structural inconsistencies).
@@
    Panel refinement. The three-expert panel reviews all feedback and updates the taxonomy, after which the next annotation round begins.
@@
    Convergence and finalization. The process terminates when experts agree that no new failure modes and rubrics revisions are needed, yielding the final taxonomy in Tab. 1.
@@
Correlation With Human Preferences. On OpenRubrics and AutoRubrics, we use judge–human preference agreement as an external validation signal for RIFT and find that rubrics with more RIFT failure mode labels are significantly more likely to exhibit judge–human misalignment (Pearson’s $r$ = 0.162, $p$ = 0.0021; average count = 3.66 vs. 3.15, $p<$ 1e-4) failure modes for misaligned vs. aligned cases), providing quantitative evidence that RIFT captures rubric properties relevant to downstream human-aligned evaluation.
@@
Table 2: Prevalence of RIFT failure modes in human-crafted vs. synthetic rubrics. Entries are the fraction of rubrics in the subset annotated with the failure mode.
Subjective Non-Atomic Ungrounded     Misaligned /Rigid Missing     Redundant Low Signal Hackable
Human-curated 52.6% 26.3% 42.1%     63.2% 47.4%     26.3% 21.1% 0.0%
Synthetic 86.7% 60.0% 46.7%     20.0% 36.7%     23.3% 40.0% 13.3%
@@ 4 Automated RIFT Evaluators
To enable scalable assessment of rubric quality under RIFT, we develop automated evaluators for each failure mode. We evaluate these signals on a held-out test set of 50 rubrics drawn from AdvancedIF, ResearchRubrics, WildChecklists, OpenRubrics, and AutoRubrics, which are manually annotated using the RIFT taxonomy. We implement several automated signals and examine how well they align with expert diagnostic labels:
  1. 1.
    LLM-as-Judge (LLMaJ). A rubric-conditioned failure-mode classifier trained using grounded-theory annotations and the RIFT taxonomy. We use GPT-5.2 as the classifier and provide the taxonomy as part of the prompt (see Appendix B).
  2. 2.
@@
Table 3: Automated RIFT evaluator alignment, as measured by F1, with expert annotations.
Failure mode LLMaJ F1 IRR F1 Alignment F1 Reward variance F1
Subjective 0.861 0.843 0.854 0.843
Non-Atomic 0.644 0.667 0.656 0.629
Ungrounded 0.444 0.625 0.625 0.717
Misaligned or Rigid 0.750 0.554 0.554 0.562
Missing Criteria 0.686 0.585 0.576 0.567
Hackable 0.000 0.400 0.400 0.267
Low Signal 0.667 0.525 0.517 0.519
Redundant Criteria 0.643 0.429 0.444 0.467
Alignment with Expert Annotations. Results are reported in Tab. 3. For each evaluator, failure-mode pair, we report F1 at the threshold that maximizes F1 on the test set. The RIFT LLMaJ achieves moderate to good alignment on a subset of failure modes, Subjective and Misaligned or Rigid. However, there is significant room for improvement of the LLMaJ on the Non-Atomic, Ungrounded, and Hackable failure modes. We recommend using non-LLMaJ-based methods to complement the LLMaJ in these cases as a concrete strategy to improve detection for these challenging categories. We expect further refinements of the LLMaJ via prompt tuning and curating in-context examples would further improve the performance of the LLMaJ.
Grouping failure modes into reliability, content, and consequential categories highlights observed difficulties of automated rubric-quality evaluation. Reliability failures are identified more accurately, likely because they result in inconsistent rubric-based judgments that non-LLMaJ signals measure directly, and because RIFT definitions provide a clear scoring guidance for LLMaJs. In contrast, content and consequential failures require detecting substance gaps or inaccuracies in rubric criteria and anticipating downstream effects.
5 Conclusions and Future Work
@@
We also implemented automated RIFT evaluators to detect and label failure modes, enabling scalable analysis. We expect expert-in-the-loop workflows to remain important for assessing content and consequential validity. Experts can verify coverage, relevance, and missing criteria (with LLM critics to accelerate annotation (McAleese et al., 2024)), while consequential validity may require methods that directly measure a rubric’s impact in downstream use cases.
RIFT is not exhaustive. Scaling the annotation process and validating RIFT more broadly (e.g., across domains, tasks, and stakeholder settings) are important directions for future work. Moreover, our there is room for improving our automated evaluators’ alignment with expert judgments. Finally, future work should explore techniques that more directly measure the downstream impacts of each failure mode and measure the impact of RIFT guided rubric refinements on training pipeline or evaluation benchmark outcomes.
Appendix A Appendix
Appendix B Prompts

### 20 Apr 2026 v2

@@ Abstract
Rubric-based evaluation is widely used in LLM benchmarks and training pipelines for open-ended, less verifiable tasks. While prior work has demonstrated the effectiveness of rubrics using downstream signals such as reinforcement learning outcomes, there remains no principled way to diagnose how a rubric itself fails from such aggregated or downstream signals alone. To address this gap, we introduce RIFT: RubrIc Failure mode Taxonomy, a taxonomy for systematically characterizing failure modes in rubric composition and design. RIFT consists of eight failure modes organized into three high-level categories: Reliability Failures, Content Validity Failures, and Consequential Validity Failures. RIFT is developed using grounded theory by iteratively annotating rubrics drawn from five diverse data sources spanning general instruction following, code generation, creative writing, and expert-level deep research, until no new failure modes are identified. We evaluate the consistency of the taxonomy by measuring agreement among independent human annotators, observing fair agreement overall (87% pairwise agreement and 0.64 average Cohen’s kappa). Finally, to support scalable diagnosis, we propose automated rubric quality metrics and show that they align with human failure-mode annotations, achieving up to 0.925 F1.
RIFT: A RubrIc Failure Mode Taxonomy and Automated Diagnostics
Zhengyang Qi Snorkel AI          Charles Dickens Snorkel AI          Derek Pham Snorkel AI          Amanda Dsouza Snorkel AI
Armin Parchami Snorkel AI          Frederic Sala Snorkel AI University of Wisconsin–Madison          Paroma Varma Snorkel AI
@@ 2 Related Work
Issues With LLM Evaluation Benchmarks. Recent studies show that LLM benchmarks suffer from unreliable outcome verification, including widely used benchmarks such as SWE-bench-Verified and $\tau$-bench (Chowdhury et al., 2024; Yao et al., 2024; Zhu et al., 2025b). Beyond outcome verification, naive LLM-as-Judge evaluation can introduce systematic bias and overconfident estimates which requires calibrated reporting (Lee et al., 2026). Prior work also reports substantial evaluation variance (Madaan et al., 2024) and other weaknesses in current benchmarking practices (Reuel et al., 2024; Eriksson et al., 2025), motivating the need for more reliable evaluation design.
Operationalization and Validity in Measurement. RIFT’s three failure categories draw on established measurement-theory concepts. Reliability Failures parallel the problem of operational definitions (Bridgman, 1927; Stevens, 1935): when rubric criteria lack precise operationalizations, different judges interpret them differently, producing inconsistent scores (Breznau et al., 2022; Carpentras, 2024). Content Validity Failures build on the psychometric notion of content validity—whether an instrument adequately represents the construct it measures (Sireci, 1998)—applied here to rubric criteria that miss or misrepresent the evaluation target. Consequential Validity Failures draw on consequential validity (Iliescu and Greiff, 2021), capturing rubrics whose design enables gaming or low discrimination downstream.
@@ 3 The RubrIc Failure mode Taxonomy
Table 1: RIFT: RubrIc Failure Mode Taxonomy. Taxonomy is developed through 4 rounds of annotations through the grounded theory by three experts. See Appendix D for complete descriptions.
@@
Reliability Failures Subjective The rubric uses inherently subjective evaluative terms (e.g., “clear,” “appropriate”) and does not anchor them with objective expectations. 0.59 0.60 86.7%
Non-Atomic The rubric does not provide a parseable, consistently scorable structure, or uses bundled criteria that prevent consistent partial credit. 0.54 0.54 86.7%
Ungrounded The rubric includes one or more criteria that require verification against information that is in principle groundable or factually checkable, but fails to provide sufficient grounding or bounds (e.g., reference facts, sources, time scope, or a verification procedure). 0.73 0.74 86.7%
Content Validity Misaligned or Rigid The rubric (a) grades the wrong objective for the prompt or embeds incorrect assumptions, or (b) imposes unnecessarily strict or narrow requirements not asked for by the prompt. 0.57 0.58 80.0%
Missing Criteria The prompt implies at least one checkable requirement, but the rubric provides no criterion that allows a grader to evaluate that requirement. 0.69 0.69 86.7%
Consequential Hackable The rubric is gameable at the rubric level: a responder could easily achieve a top score by inflating proxy metrics without materially improving correctness or fulfillment of the prompt. 0.74 0.72 93.3%
Low Signal The rubric as a whole does not discriminate candidate responses well, i.e., it would give nearly equivalent scores to many substantively different-quality responses. 0.86 0.86 93.3%
Redundant Criteria Two or more rubric criteria evaluate the same underlying requirement, such that the same behavior is rewarded or penalized multiple times. 0.70 0.67 86.7%
@@
We define the RubrIc Failure mode Taxonomy (RIFT) as a generic framework for identifying and classifying failure modes in evaluation rubrics (Tab. 1). To avoid imposing a top-down schema, RIFT is developed using grounded theory (Glaser and Strauss, 1967a), in which failure modes are iteratively derived from expert annotations and open-ended feedback. The resulting taxonomy organizes rubric failures into three core categories: Reliability (consistency and reproducibility of judgments; cf. the literature on operational definitions (Bridgman, 1927; Stevens, 1935)), Content Validity (alignment between rubric criteria and the intended evaluation target (Sireci, 1998)), and Consequential Validity (downstream usefulness and discriminative power of the rubric (Iliescu and Greiff, 2021)).
@@
Taxonomy Development with Grounded Theory. We develop RIFT through an iterative process inspired by Grounded Theory, following prior work on failure taxonomies for multi-agent LLM systems (Cemri et al., 2025) and practitioner guide in (Chun Tie et al., 2019). The process spans four sequential annotation rounds and covers 85 rubrics in total. Ae use purposive sampling to select rubrics from AdvancedIF, ResearchRubrics, WildChecklists, OpenRubrics, and AutoRubrics. These sources were chosen to maximize variation in rubric origin (expert-written or auto-generated), format (checklists, principles, or narrative rubrics), length, and domain. Full details on data composition are provided in Appendix A.
Three experts develop the taxonomy iteratively using standard Grounded Theory practices (Chun Tie et al., 2019). We use open coding to derive candidate failure modes from open-ended rubric critiques, constant comparative analysis to compare new observations against existing categories, and memoing to record emerging failure modes, boundary cases, and revision decisions. Data collection and analysis proceed concurrently across rounds: each annotation round informs the next revision of the taxonomy. The process continues until saturation condition is reached, which we define as the point at which additional rubrics no longer reveal new failure modes or require substantive taxonomy revisions. We operationalize this process as follows:
@@
    Bootstrapping (first round only). Experts independently write open-ended rubric critiques for each rubric. These critiques are provided to GPT-5.2 to propose an initial candidate taxonomy.
@@
    Labeling and feedback. Experts label each rubric using the current taxonomy. They also document failure modes not covered by existing categories and note issues in the taxonomy itself, such as unclear definitions, overlapping categories, or structural inconsistencies.
@@
    Panel refinement. The three-expert panel reviews all labels, critiques, and feedback, then revises the taxonomy before the next annotation round.
@@
    Convergence and finalization. The process terminates when experts agree that no new failure modes are emerging and no further rubric-driven taxonomy revisions are needed, yielding the final taxonomy in Tab. 1.
A full per-source, per-round provenance breakdown is provided in Appendix A.
@@
Correlation With Human Preferences. Among the five data sources, only OpenRubrics and AutoRubrics provide paired judge-human preference labels, so we restrict this analysis to those two sources. We use judge-human preference agreement as an external validation signal for RIFT and find that rubrics with more RIFT failure mode labels are significantly more likely to exhibit judge–human misalignment (Pearson’s $r$ = 0.162, $p$ = 0.0021; average count = 3.66 vs. 3.15, $p<$ 1e-4; effective size = 20) failure modes for misaligned vs. aligned cases), providing quantitative evidence that RIFT captures rubric properties relevant to downstream human-aligned evaluation.
@@
Table 2: Prevalence of RIFT failure modes in human-crafted vs. synthetic rubrics. Entries are the fraction of rubrics in each subset annotated with the failure mode. Category averages are the mean prevalence across failure modes within each category.
Failure mode Human Synthetic
Subjective 52.6% 86.7%
Non-Atomic 26.3% 60.0%
Ungrounded 42.1% 46.7%
Reliability Avg. 40.3% 64.5%
Misaligned/Rigid 63.2% 20.0%
Missing 47.4% 36.7%
Content Validity Avg. 55.3% 28.4%
Redundant 26.3% 23.3%
Low Signal 21.1% 40.0%
Hackable 0.0% 13.3%
Consequential Validity Avg. 15.8% 25.5%
@@ 4 Automated RIFT Evaluators
Table 3: Automated RIFT evaluator alignment, as measured by F1, with expert annotations. LLMaJ columns report single-run and majority-vote (N=5N{=}5) F1 for GPT-5.2 and Gemini 3 Pro as the classifier
Failure mode GPT-5.2 LLMaJ Gemini 3 Pro LLMaJ IRR F1 Align F1 Reward var. F1
Single MV Single MV
Subjective 0.861 0.892 0.909 0.925 0.843 0.854 0.843
Non-Atomic 0.644 0.702 0.630 0.679 0.667 0.656 0.629
Ungrounded 0.444 0.529 0.634 0.651 0.625 0.625 0.717
Misaligned or Rigid 0.750 0.778 0.737 0.718 0.554 0.554 0.562
Missing Criteria 0.686 0.706 0.452 0.483 0.585 0.576 0.567
Hackable 0.000 0.000 0.000 0.000 0.400 0.400 0.267
Low Signal 0.667 0.625 0.400 0.316 0.525 0.517 0.519
Redundant Criteria 0.643 0.500 0.417 0.435 0.429 0.444 0.467
Table 4: Pairwise agreement between majority-vote LLMaJ models reported per failure mode and macro-averaged (Overall). All models use majority-vote aggregation over N=5N{=}5 independent runs.
Failure mode GPT-5.2 $\leftrightarrow$ Gemini 3 Pro GPT-5.2 $\leftrightarrow$ Sonnet 4.5 Sonnet 4.5 $\leftrightarrow$ Gemini 3 Pro
% Agree $\kappa$ % Agree $\kappa$ % Agree $\kappa$
Subjective 0.820 0.573 0.860 0.471 0.720 0.268
Non-Atomic 0.720 0.367 0.800 0.453 0.640 0.168
Ungrounded 0.640 0.240 0.800 0.593 0.560 0.116
Misaligned or Rigid 0.780 0.547 0.700 0.266 0.600 0.120
Missing Criteria 0.820 0.549 0.780 0.503 0.720 0.318
Hackable 0.920 0.000 0.960 0.779 0.880 0.000
Low Signal 0.720 0.220 0.820 0.540 0.900 0.502
Redundant Criteria 0.720 0.234 0.780 0.347 0.820 0.419
Overall (macro) 0.768 0.341 0.812 0.494 0.730 0.239
To enable scalable assessment of rubric quality under RIFT, we develop automated evaluators for each failure mode. We evaluate these signals on an independent test set of 50 rubrics (10 per source, stratified equally across all five data sources) sampled without replacement from rubrics not used during taxonomy development and manually annotated using the RIFT taxonomy. All 50 rubrics are evaluated by every automated signal; for both LLMaJ classifiers (GPT-5.2 and Gemini 3 Pro), each rubric is evaluated by both models. We implement several automated signals and examine how well they align with expert diagnostic labels:
  • LLM-as-Judge (LLMaJ). A rubric-conditioned failure-mode classifier trained using grounded-theory annotations and the RIFT taxonomy. We report results for two classifiers, GPT-5.2 and Gemini 3 Pro, and provide the taxonomy as part of the prompt (see Appendix C).
  • Majority-Vote LLMaJ (LLMaJ MV). The same rubric-conditioned classifiers run N=5N{=}5 times independently per rubric. Per-failure-mode labels are aggregated by strict majority vote: a label is retained only if at least $\lceil N/2\rceil$ runs predict it.
@@
Alignment with Expert Annotations. Results are reported in Tab. 3. For each evaluator, failure-mode pair, we report F1 at the threshold that maximizes F1 on the test set. The single-run LLMaJs achieves moderate to good alignment on a subset of failure modes, Subjective and Misaligned or Rigid, but exhibits notable misalignment on Non-Atomic and Hackable. Aggregating N=5N{=}5 independent LLMaJ runs via majority vote (LLMaJ MV) typically improves performance on the eight failure modes over the single-run LLMaJ and becomes the best evaluator on four. Majority voting degrades Low Signal and Redundant Criteria, where the non-LLMaJ signals and single-run LLMaJ, respectively, remain the strongest evaluators. GPT-5.2 and Gemini 3 Pro have noticeably different strengths: Gemini 3 Pro is substantially better on Ungrounded (0.634 vs. GPT-5.2’s 0.444) and Subjective (0.909 vs. 0.861), but weaker on Missing Criteria (0.452 vs. 0.686) and Low Signal (0.400 vs. 0.667), indicating that different frontier models surface complementary failure modes. The non-LLMaJ-based methods complement the LLMaJ-based evaluators, and we expect further gains from prompt tuning, curating in-context examples, and combining majority voting with these complementary signals.
Grouping failure modes into reliability, content, and consequential categories highlights observed difficulties of automated rubric-quality evaluation. Reliability failures are identified more accurately, likely because they result in inconsistent rubric-based judgments that non-LLMaJ signals measure directly, and because RIFT definitions provide clear scoring guidance for LLMaJs. In contrast, content and consequential failures require detecting substance gaps or inaccuracies in rubric criteria and anticipating downstream effects.
Pairwise Model Agreement. Tab. 4 reports per failure mode and macro-averaged agreement between three frontier majority-vote LLMaJs (Sonnet 4.5, Gemini 3 Pro, GPT-5.2) on the 50-rubric RIFT test set. Overall cross-model agreement rates range from $0.730$ to $0.812$ and Cohen’s $\kappa$ from $0.239$ to $0.494$. Between Gemini 3 Pro and GPT-5.2, per failure mode, Cohen’s $\kappa$ ranges from $0.0$ to $0.573$. This suggests moderate taxonomy interpretation differences between current frontier models and motivates leveraging experts in the loop for RIFT evaluations on failure modes such as Consequential failure modes with low model agreement.
5 Conclusions
@@
We also implemented automated RIFT evaluators to detect and label failure modes, enabling scalable analysis. In practice, we envision RIFT supporting an iterative rubric improvement workflow, for instance: (1) diagnose: run automated evaluators to flag likely failure modes for a candidate rubric, (2) review: an expert inspects flagged failures using the taxonomy’s decision rules and confirms or dismisses each, (3) revise: apply targeted fixes guided by the failure mode (e.g., anchor subjective terms with concrete criteria, unbundle non-atomic items into separately scored sub-criteria, or add missing requirements), and (4) validate: re-run the automated signals to verify the revision resolved the flagged issues. We expect expert-in-the-loop workflows to remain important for assessing content and consequential validity. Experts can verify coverage, relevance, and missing criteria (with LLM critics to accelerate annotation (McAleese et al., 2024)), while consequential validity may require methods that directly measure a rubric’s impact in downstream use cases.
6 Limitations
RIFT is not intended to be an exhaustive taxonomy of rubric failure modes. The taxonomy is derived from the rubrics, artifacts, and annotation settings examined in this study, and additional failure modes may emerge in other domains, task formats, or stakeholder contexts. As a result, the current taxonomy should be viewed as a useful but incomplete characterization of rubric breakdown.
The generalizability of RIFT remains only partially validated. Our annotations were conducted on a bounded set of tasks and artifacts, and rubric quality may depend heavily on the application setting, evaluation goal, and stakeholder priorities. Future work should test the taxonomy across a broader range of domains, annotator populations, and use cases.
The annotation process itself is expensive and difficult to scale. Although expert annotation provides high-quality supervision, it is still subject to ambiguity and subjectivity, particularly for abstract rubric properties such as completeness or specificity. This limits both the scale of our study and the reliability with which fine-grained failure modes can be distinguished.
Finally, evaluator alignment with expert judgments still has substantial room for improvement. Agreement with expert labels is only an indirect proxy for practical utility, and future work should more directly measure how identifying different failure modes affects downstream outcomes such as rubric revision quality, evaluator robustness, and benchmark reliability.
7 Ethical Statement
Rubrics are often used in high-impact evaluation settings, so errors in rubric design can propagate into misleading judgments, poor model selection, or unfair downstream decisions. Our goal in introducing RIFT is to support more transparent and reliable rubric development by making common failure modes easier to identify and analyze. At the same time, we do not view automated rubric diagnosis as a substitute for human oversight, especially in sensitive or high-stakes applications. We believe expert review remains essential for ensuring that evaluation criteria are appropriate, fair, and aligned with the intended use case.
Appendix A Data Source and Sampling Provenance
This appendix summarizes the data sources and sampling procedure used for taxonomy development (Section 3) and automated evaluator experiments (Section 4).
A.1 Data source distribution.
We sample rubrics from five datasets selected to cover variation in rubric origin, format, length, and task domain. Tab. 5 summarizes each source.
Table 5: Summary of rubric data sources. Top domains are listed in descending frequency within the sampled rubrics from each dataset.
Dataset Origin Format Mean Words Top Domains
AutoRubrics LLM-generated Narrative 68 Coding, General, Creative
WildChecklists LLM-generated Checklist 111 General, Creative, Coding
OpenRubrics LLM-generated Principles 145 General, Coding, Health
ResearchRubrics Expert-written Checklist 1,059 Creative, Health, Coding, Education
AdvancedIF Expert-curated Checklist 88 Coding, Creative, General, Education
A.2 Sampling.
Rubrics are drawn by stratified uniform random sampling with equal counts per source, without replacement across rounds, and with a fixed random seed per round for reproducibility. The taxonomy development sample contains 85 rubrics across four rounds: Round 1 collects open-ended critiques for bootstrapping, while Rounds 2–4 use the evolving taxonomy for labeling. Each rubric is independently annotated by at least three experts. The test set is drawn independently with the same stratified procedure, contains 50 rubrics with no overlap with the taxonomy development sample, and is evaluated by all automated signals in Section 4. For the LLM-as-Judge classifiers, both GPT-5.2 and Gemini 3 Pro evaluate every test rubric. Tab. 6 summarizes both samples.
Table 6: Per-source rubric counts for taxonomy development and automated-evaluator testing.
Source Rnd 1 Rnd 2 Rnd 3 Rnd 4 Test
AdvancedIF 5 5 5 2 10
ResearchRubrics 5 5 5 2 10
WildChecklists 5 5 5 2 10
OpenRubrics 5 5 5 2 10
AutoRubrics 5 5 5 2 10
Total 25 25 25 10 50
Appendix B Additional Experiment Details
This appendix provides additional analyses for the automated RIFT evaluators. First, we test whether the Alignment and Reward Variance signals are sensitive to the choice of judge model by replacing GPT-5.2 with Gemini 3 Pro Preview. Second, we report threshold-free ROC-AUC for the non-LLMaJ evaluators to complement best-threshold F1 and better account for class imbalance in the failure-mode labels.
B.1 Alignment and Reward Variance Judge Model Ablation
Table 7: Judge-model ablations for Alignment and reward variance.
Failure mode Alignment (F1) Reward variance (F1)
GPT-5.2 ref. Gemini 3 Pro Preview ref. GPT-5.2 judge Gemini 3 Pro Preview judge
Subjective 0.854 0.854 0.843 0.843
Non-Atomic 0.656 0.647 0.629 0.629
Ungrounded 0.625 0.646 0.717 0.618
Misaligned or Rigid 0.554 0.567 0.562 0.567
Missing Criteria 0.576 0.603 0.567 0.567
Hackable 0.400 0.276 0.267 0.400
Low Signal 0.517 0.508 0.519 0.558
Redundant Criteria 0.444 0.436 0.467 0.455
The main paper reports two automated signals whose “judge” role is instantiated with GPT-5.2. Alignment measures how often weaker preference labelers (GPT-5 mini, Claude Haiku 4.5, and Gemini 3 Flash) agree with a strong reference labeler (GPT-5.2) on the same rubric-conditioned response pairs. We ablate by swapping the reference labeler to Gemini 3 Pro Preview, holding the three weaker labelers and the response pairs fixed. Reward variance measures stability of the aggregate rubric score under a rubric-conditioned LLM-as-Judge (LLMaJ) that scores four independent responses from GPT-5 mini per test input; we ablate by swapping that judge from GPT-5.2 to Gemini 3 Pro Preview, holding the response generator and evaluation protocol fixed. As shown in Tab. 7, both signals are largely judge-model robust on Subjective and Non-Atomic, with the largest swings on consequential failure modes (notably Hackable and Low Signal). We hypothesize that is caused by label imbalance in the evaluation dataset.
B.2 Additional Analysis for non-LLMaJ RIFT Evaluators
We additionally report threshold-free ROC-AUC results for automated failure-mode detection to address the concern that F1 at the best threshold on the same evaluation set may overstate performance due to in-sample threshold tuning. In particular, some failure modes are highly imbalanced (e.g., only 4 positives for Hackable), making threshold-dependent F1 especially unstable. We therefore compute direction-agnostic ROC-AUC as $\max(\mathrm{AUC},1-\mathrm{AUC})$, matching the direction-agnostic threshold sweep used for F1.
We evaluate all evaluators except LLMaJ, including IRR, Alignment, Reward variance. Across the 48 joined rubrics, IRR has mean 0.712 (std. 0.149), Alignment has mean 0.704 (std. 0.166), and Reward Variance has mean 0.025 (std. 0.053).
Tables 8 and 9 show that threshold-free AUC is generally more conservative than best-threshold F1. The clearest example is Subjective, where F1 reaches 0.854 but best AUC is only 0.640, suggesting that much of the F1 comes from class prevalence rather than strong ranking ability. In contrast, Hackable has the strongest AUC (0.807 with Alignment) despite a modest best F1 of 0.400, indicating unstable signals under severe class imbalance. The strongest diagnostic also differs by failure mode, i.e., Alignment works best for Hackable, Misaligned or Rigid, and Missing Criteria; IRR works best for Non-Atomic; and Reward Variance works best for Ungrounded, Low Signal, and Redundant Criteria, suggesting that the three signals provide complementary information. Finally, correlations with the total number of failure modes are weak ($r=0.247$ for IRR, $r=0.222$ for Alignment, and $r=0.052$ for Reward Variance), so these signals are more informative about which failure mode is present than how many are present overall.
Table 8: Threshold-free ROC-AUC and best-threshold F1 for automated failure-mode detection. AUC is computed direction-agnostically as $\max(\mathrm{AUC},1-\mathrm{AUC})$.
Failure mode $n_{+}$ ROC-AUC Best F1
IRR Align Var IRR Align Var
Subjective 35 0.547 0.545 0.640 0.843 0.854 0.843
Non-Atomic 22 0.581 0.512 0.506 0.667 0.656 0.629
Ungrounded 21 0.532 0.502 0.644 0.625 0.625 0.717
Misaligned or Rigid 18 0.508 0.537 0.532 0.554 0.554 0.563
Missing Criteria 19 0.541 0.599 0.532 0.585 0.576 0.567
Hackable 4 0.759 0.807 0.656 0.400 0.400 0.267
Low Signal 16 0.515 0.537 0.543 0.525 0.517 0.519
Redundant Criteria 12 0.507 0.517 0.556 0.429 0.444 0.467
Table 9: Comparison between the best threshold-free AUC and best in-sample F1 across diagnostic signals.
Failure mode Best signal (AUC) Best AUC Best F1
Subjective Reward Variance 0.640 0.854
Non-Atomic IRR 0.581 0.667
Ungrounded Reward Variance 0.644 0.717
Misaligned or Rigid Alignment 0.537 0.563
Missing Criteria Alignment 0.599 0.585
Hackable Alignment 0.807 0.400
Low Signal Reward Variance 0.543 0.525
Redundant Criteria Reward Variance 0.556 0.467
Appendix C Prompts